# M2IST: Multi-Modal Interactive Side-Tuning for Memory-efficient Referring Expression Comprehension

### v1

@@ 3.2 Multi-Modal Interactive Side-Tuning
  • Fully freezing the pre-trained encoders. We choose to directly keep the pre-trained parameters fixed and only fine-tune the V-L Encoder. While it effectively saves a significant amount of GPU memory, it also results in significantly inferior performance (see Table 3 (a)).
  • Updating a few additional parameters. We explore various mainstream PETL methods. Though most of them achieve relatively satisfactory performance as well as save tunable parameters, updating the additional parameters still necessitates substantial GPU memory rather than effectively mitigating the computational load (see Table 2).
Besides, it is obvious that only the V-L encoder is responsible for cross-modality fusion. However, such fusion does not exist in the shallow layers of the base model, which is insufficient when the given referring sentence contains complex semantic information, such as spatial relations Huang and Satoh (2023). To address the above issues, we propose Multi-Modal Interactive Side-Tuning (M2IST) that keeps the pre-trained encoders frozen and updates the proposed Mixture of Multi-Modal Interactive Side Adapters (M3ISA) on side networks to facilitate parameter- and memory-efficient fine-tuning for REC, as shown in Figure 2. Note that we do not show the LayerNorm for simplicity.
M3ISA architecture.
The core component of M2IST is M3ISA (see Figure 2 (right)), which consists of two distinct adapters (intra- and inter-modality adapters) to effectively and efficiently bridge the Vision Encoder and Language Encoder. The intra-modality adapters follow the basic design of Adapter Houlsby et al. (2019) in NLP, and include Vision Expert Adapter (VEA) and Language Expert Adapter (LEA), shown as separate blue branch and green branch in Figure 2 (right). Both of them consist of a down-projection layer $\mathbf{W}_{\text{down}}$, ReLU non-linear activation, and an up-projection layer $\mathbf{W}_{\text{up}}$ in sequence. They are responsible for transferring the pre-trained single-modality representations to more fine-grained ones for the REC domain. Specifically, taking the VEA as an example, given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, the function of VEA can be formally expressed as:
@@
The inter-modality adapters, Interaction Expert Adapters (IEA) are designed to enhance cross-modality interactions by progressively bridging the pre-trained dual encoders, inspired by existing efforts Zhou and Long (2023a); Xu et al. (2023). As depicted by the entire pink section in Figure 2 (right), IEA include a unique down-projection layer for vision $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.180,0.459,0.714}C_{v}}\times C_{d}}$ and language $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.439,0.678,0.278}C_{l}}\times C_{d}}$, ReLU activation, an interactive up-projection layer $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times{\color[rgb]{0.929,0.420,0.380}C_{i}}}$, and a unique up-projection layer for vision $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.180,0.459,0.714}C_{v}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$ and language $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.439,0.678,0.278}C_{l}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$, where $C_{v}$, $C_{l}$, and $C_{i}$ represent the vision, language, and interaction channels, respectively. Given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and language tokens $\bm{x}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$, the corresponding down-projection layers first down-sample them to the bottleneck features $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{d}}$ and $\bm{z}_{l}\in\mathbb{R}^{N_{l}\times C_{d}}$. Then, the corresponding up-projection layers and interactive up-projection layer up-sample these bottleneck features and concatenate them within the same modality to obtain the cross-modality features $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$ as:
@@
As depicted in Figure 2 (left), we incorporate a stack of M3ISAs into two side networks that operate in parallel with the pre-trained dual encoders. Specifically, in one encoder layer (both for vision and language), the IEA first receives processed vision/language tokens from the Multi-head Attention (MHA) layers as input and produces adapted, interacted tokens for the vision/language side network. Subsequently, the VEA/LEA take the processed vision/language tokens from the Feed Forward Networks (FFN) as input and generate adapted single-modality tokens for the corresponding side networks. The outputs of the IEA and VEA/LEA are added within the vision/language side networks, along with the original vision/language tokens through skip-connections. After passing through the side networks, the outputs of the vision/language side networks are added to the outputs of the vision/language encoders. During fine-tuning, we keep the pre-trained vision and language encoders fixed and update the M3ISAs in the side networks, allowing the pre-trained encoders to act as standalone feature extractors. Pseudocode is presented in Appendix G.
@@
VC Zhang et al. (2018) VGG16 LSTM 17 - 73.33 67.44 - 58.40 53.18 62.30 - -
ParalAttn Zhuang et al. (2018) VGG16 LSTM 17 - 75.31 65.52 - 61.34 50.86 58.03 - -
MAttNet Yu et al. (2018) RN101 LSTM 47 76.65 81.14 69.99 65.33 71.62 56.00 - 66.58 67.27
RvG-Tree Hong et al. (2019) RN101 LSTM 47 75.06 78.61 69.85 63.51 67.45 56.66 - 66.95 66.51
@@
FAOA Yang et al. (2019) DN53 LSTM 43 72.54 74.35 68.50 56.81 60.23 49.60 56.12 61.33 60.26
RCCF Liao et al. (2020b) DLA34 LSTM 18 - 81.06 71.85 - 70.35 56.32 - - 65.73
ReSC Yang et al. (2020) DN53 BERT 152 76.59 78.22 73.25 63.23 66.64 55.53 63.12 67.30 67.20
RealGIN Zhou et al. (2021) DN53 GRU 41 77.25 78.70 72.10 62.78 67.17 54.21 - 62.75 62.33
TransVG Deng et al. (2021) RN50 BERT 151 80.49 83.28 75.24 63.50 68.15 55.63 66.56 67.66 67.44
VGTR Du et al. (2022) RN50 LSTM 52 78.70 82.09 73.31 63.57 69.65 55.33 62.88 65.62 65.30
PFOS Sun et al. (2022) DN53 BERT 152 77.37 80.43 72.87 63.74 68.54 55.84 61.46 67.08 66.35
SeqTR Zhu et al. (2022) DN53 GRU 41 78.22 81.47 73.80 66.01 70.23 55.68 - 68.26 -
DMRNet Zhang et al. (2023) DN53 BERT 152 76.99 79.71 72.67 61.58 66.60 54.00 - 66.03 66.70
M2IST (Ours) RN50 BERT 3.19 81.35 82.29 77.98 63.15 67.11 55.52 67.50 67.67 67.41
Table 1: Comparison with full fine-tuning on RefCOCO, RefCOCO+, and RefCOCOg. "RN50", "RN101", and "DN53" represent ResNet-50, ResNet-101, and DarkNet-53 respectively. and denote RN50 followed by 6 and 2 transformer encoder layers respectively. "Param." shows the number of tunable encoder parameters.
(M) (GB) val testA testB val testA testB val-g val-u test-u
Fully fine-tuning 151 38.95 80.49 83.28 75.24 63.50 68.15 55.63 66.56 67.66 67.44
Adapter Houlsby et al. (2019) 3.27 28.52 78.02 79.89 75.23 61.35 66.34 54.21 63.18 65.26 66.65
LoRA Hu et al. (2022) 2.37 20.37 77.57 78.22 73.37 61.24 66.53 53.95 64.27 67.36 66.43
AdaptFormer Chen et al. (2022) 2.38 20.37 76.32 77.16 73.94 60.96 65.19 53.88 61.81 65.44 64.37
CM Adapter Jiang et al. (2022) 3.27 27.19 77.37 78.81 74.07 61.34 66.10 53.31 63.93 65.75 64.72
MRS-Adapter Yuan et al. (2023) 1.58 20.07 77.14 77.80 74.80 61.13 66.38 53.13 63.07 66.46 65.16
M2IST (Ours) 3.19 15.44 81.35 82.29 77.98 63.15 67.11 55.52 67.50 67.67 67.41
Table 2: Comparison with PETL methods using the same base architecture on RefCOCO, RefCOCO+ and RefCOCOg. "Param." indicates the number of tunable parameters in the pre-trained encoders. "Mem." denotes the peak GPU memory footprint with batch size 64 during fine-tuning.
3.3 Discussion: Advantages of M2IST
The proposed M2IST offers several advantages over fully fine-tuning and other PETL methods, summarized as follows:
Parameter Efficiency. Fully fine-tuning pre-trained encoders is computationally expensive due to their large size and complexity Liu et al. (2024c). Furthermore, it often leads to forgetting valuable pre-trained knowledge and increases the risk of overfitting, as the encoders are fine-tuned on limited data. M2IST mitigates these issues by freezing the pre-trained encoders and updating only the lightweight M3ISAs, achieving effective intra- and inter-modality representation adaptation and enhanced performance (see Table 3 (f)).
Memory Efficiency. Both full fine-tuning and other PETL methods require backpropagation through large pre-trained encoders, leading to high GPU memory usage. M2IST reduces this by separating tunable parameters from the pre-trained encoders and placing them in parallel side interactive networks. These networks facilitate single-modality knowledge transfer and enable progressive cross-modality interaction, enhancing deep vision-language alignment by the V-L Encoder. Since gradients backpropagate through the lightweight M3ISAs instead of the heavy encoders, GPU memory requirements are significantly reduced. Additionally, M2IST maintains the baseline model’s architecture, simplifying its implementation compared to other PETL methods.
4 Experiments
4.1 Experimental Setup
Datasets and Evaluation Metrics. We conduct experiments on the widely-used REC benchmarks: RefCOCO Yu et al. (2016), RefCOCO+ Yu et al. (2016), and RefCOCOg Mao et al. (2016); Nagaraja et al. (2016). More dataset details are provided in Appendix A. We use Precision@0.5 as the evaluation metric. In addition to accuracy, we also report the number of tunable parameters in the pre-trained encoders and the training memory consumption in Gigabytes (GB) to compare the fine-tuning efficiency with other PETL methods.
Implementation Details. The Vision Encoder is initialized with ResNet-50 He et al. (2016) and the DETR encoder Carion et al. (2020), while the Language Encoder is initialized with BERT-base Devlin et al. (2018). The bottleneck dimension $C_{d}$ for VEA/LEA is 128, and the interaction dimension $C_{i}$ for IEA is 256. For fair comparisons, all PETL methods use the same base architecture, keeping the Vision and Language Encoders fixed while updating only the V-L Encoder during fine-tuning. More details are provided in Appendix B.
4.2 Main Results
Table 1 demonstrates that M2IST achieves competitive performance across three benchmarks compared to full fine-tuning methods. Specifically, on the three sets of RefCOCOg Mao et al. (2016); Nagaraja et al. (2016), M2IST outperforms the majority of other baseline methods. Two-stage REC methods achieve outstanding performance in RefCOCO+ Yu et al. (2016) because the referring sentences in the RefCOCO+ dataset only describe the appearance and attributes of objects. Two-stage REC methods can more explicitly locate referred objects by directly computing the similarity scores between region proposals and sentences. Even so, Table 1 illustrates that M2IST achieves an optimal performance-parameter trade-off compared to full fine-tuning methods, underscoring its advantage in parameter efficiency, as discussed in Section 3.3.
@@ 4.2 Main Results
Table 2 illustrates that M2IST outperforms other PETL methods on all three benchmarks. This highlights the effectiveness of M3ISAs in adapting pre-trained knowledge for the REC domain. Furthermore, through the facilitation of cross-modality interaction between the encoders, M3ISAs enhance the modeling of complex spatial relationships, leading to improved performance on RefCOCOg Mao et al. (2016); Nagaraja et al. (2016). Regarding fine-tuning efficiency, M2IST requires the least training memory among PETL methods. This results from the fact that gradients backpropagate through the lightweight M3ISAs rather than the heavy encoders, highlighting M2IST’s advantage in memory efficiency, as mentioned in Section 3.3.

### v2

@@ III-B Multi-Modal Interactive Side-Tuning: M2IST
  1. 1.
    Fully freezing the pre-trained encoders. We choose to directly keep the pre-trained parameters fixed and only fine-tune the V-L Encoder. While it effectively saves a significant amount of GPU memory, it also results in significantly inferior performance (see Table V (a)).
  2. 2.
    Updating a few additional parameters. We explore various mainstream PETL methods, such as Adapter [12] and LoRA [13]. Though most of them achieve relatively satisfactory performance as well as save tunable parameters, updating the additional parameters still necessitates substantial GPU memory rather than effectively mitigating the computational load (see Table I).
To address the above issues, we propose Multi-Modal Interactive Side-Tuning (M2IST) that keeps the pre-trained encoders frozen and updates the proposed Mixture of Multi-Modal Interactive Side Adapters (M3ISA) on side networks to facilitate parameter- and memory-efficient fine-tuning for REC, as shown in Figure 2. Note that we do not show the LayerNorm for simplicity.
M3ISA Architecture
The core component of our M2IST is M3ISA (see Figure 2 (right)), which consists of two distinct adapters, i.e., intra- and inter-modality adapters to effectively and efficiently bridge the pre-trained Vision Encoder and Language Encoder.
The intra-modality adapters include Vision Expert Adapter (VEA) and Language Expert Adapter (LEA), shown as separate blue branch and green branch in Figure 2 (right). They adhere to a fundamental design [12] for transferring pre-trained single-modality representations to more domain-specific ones. Specifically, both of them consist of a down-projection layer $\mathbf{W}_{\text{down}}$, ReLU non-linear activation, and an up-projection layer $\mathbf{W}_{\text{up}}$ in sequence. Taking the VEA as a formulaic example, given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, the function of VEA can be formally expressed as:
@@
In the base architecture, the Vision Encoder and Language Encoder are responsible for extracting single-modality features, while only the V-L encoder is tasked with cross-modality fusion through token-level vision-language alignment. However, such the late fusion has been proved insufficient when the referring sentence contains complex semantic information, such as spatial relationships [47]. Therefore, we introduce an inter-modality adapter, named the Interaction Expert Adapter (IEA). Specifically, each IEA shares a set of tokens within the same channel dimensions $C_{i}$ (i.e., Interacted Tokens in Figure 2 (right)) between the Vision Encoder and Language Encoder, thereby achieving channel-level vision-language alignment during the multi-modal feature extraction stage. As depicted by the entire pink section in Figure 2 (right), IEA include a unique down-projection layer for vision $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.180,0.459,0.714}C_{v}}\times C_{d}}$ and language $\mathbf{W}_{\text{down}}$ $\in\mathbb{R}^{{\color[rgb]{0.439,0.678,0.278}C_{l}}\times C_{d}}$, ReLU activation, an interactive up-projection layer $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times{\color[rgb]{0.929,0.420,0.380}C_{i}}}$, and a unique up-projection layer for vision $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.180,0.459,0.714}C_{v}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$ and language $\mathbf{W}_{\text{up}}$ $\in\mathbb{R}^{C_{d}\times({\color[rgb]{0.439,0.678,0.278}C_{l}}-{\color[rgb]{0.929,0.420,0.380}C_{i}})}$, where $C_{v}$, $C_{l}$, and $C_{i}$ represent the vision, language, and interaction channels, respectively. Given the vision tokens $\bm{x}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and language tokens $\bm{x}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$, the corresponding down-projection layers first down-sample them to the bottleneck features $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{d}}$ and $\bm{z}_{l}\in\mathbb{R}^{N_{l}\times C_{d}}$. Then, the corresponding up-projection layers and interactive up-projection layer up-sample these bottleneck features and concatenate them within the same modality to obtain the cross-modality features $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times{\color[rgb]{0.180,0.459,0.714}C_{v}}}$ and $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times{\color[rgb]{0.439,0.678,0.278}C_{l}}}$ as:
@@
Our IEA enables lightweight adaptation of cross-modality representations without increasing the explicit computational burden of vision-language feature extraction, thereby efficiently bridging the pre-trained vision and language encoders. Our experimental results in Table IX further demonstrate that deeper channel-level vision-language alignment yields better performance for REC. Moreover, such alignment will further improve the token-level vision-language alignment in V-L Encoder, leading to better region-sentence understanding in challenging REC scenarios, as analyzed in Section IV-E.
M3ISA Implementation
@@
As depicted in Figure 2 (left), we incorporate a stack of M3ISAs into two side networks that operate in parallel with the pre-trained dual encoders. Specifically, in one encoder layer (both for vision and language), the IEA first receives processed vision/language tokens from the Multi-head Attention (MHA) layers as input and produces adapted, interacted tokens for the vision/language side network. Subsequently, the VEA/LEA take the processed vision/language tokens from the Feed Forward Networks (FFN) as input and generate adapted single-modality tokens for the corresponding side networks. The outputs of the IEA and VEA/LEA are added within the vision/language side networks, along with the original vision/language tokens through skip-connections. After passing through the side networks, the outputs of the vision/language side networks are added to the outputs of the vision/language encoders. During fine-tuning, we keep the pre-trained encoders fixed and update the M3ISAs in the side networks, allowing the pre-trained encoders to act as standalone feature extractors.
Training Objectiveness
Following most transformer-based REC methods [3, 26, 9], the training loss function is a combination of the widely used smooth L1 loss and GIoU loss. Specifically, the prediction is donated as $\mathbf{b}=(x,y,w,h)$, and the normalized ground-truth box as $\hat{\mathbf{b}}=(\hat{x},\hat{y},\hat{w},\hat{h})$. The training objective is:
III-C Discussion: Advantages of M2IST
The proposed M2IST offers several advantages over fully fine-tuning and other PETL methods for REC, which we can summarize as three efficiency factors:
  1. 1.
    Parameter Efficiency. Fully fine-tuning pre-trained vision-language foundation models is computationally expensive due to their large size and complexity [5, 48]. Furthermore, it often leads to forgetting valuable pre-trained knowledge and increases the risk of overfitting, as the encoders are fine-tuned on limited data. M2IST mitigates these issues by freezing the pre-trained encoders and updating only the lightweight M3ISAs, achieving effective intra- and inter-modality representation adaptation and enhanced performance.
  2. 2.
    Memory Efficiency. Both full fine-tuning and other PETL methods require backpropagation through large pre-trained vision-language foundation models, leading to high GPU memory usage [43, 44]. M2IST reduces this by separating tunable parameters from the pre-trained encoders and placing them in parallel side interactive networks. Since gradients backpropagate through the lightweight M3ISAs instead of the heavy encoders, where we only need to store the parameters of our side networks, this leads to reduced GPU memory requirements. Additionally, M2IST maintains the baseline model’s architecture, simplifying its implementation compared to other PETL methods.
  3. 3.
    Time Efficiency. Most PETL methods are able to enhance the speed of fine-tuning compared to full fine-tuning, primarily due to the reduced number of updated parameters. Compared to other PETL methods, our M2IST may offer greater both fine-tuning and inference time efficiency. As for the fine-tuning stage, M2IST introduces tunable parameters in parallel with the pre-trained encoders, allowing gradient computation to occur in the lightweight M3ISAs instead of the computationally intensive encoders. As for the inference stage, compared to most PETL methods that introduce new parameters into the pre-trained networks, compromising inference efficiency, our M2IST can process the additional parameters and the pre-trained networks in parallel, making the inference more efficient.
Therefore, our approach offers parameter-, memory-, and time-efficient adaptation for visual-language foundational models, and enhances the REC task with more comprehensive visual-language alignment.
TABLE I: Comparison with PETL methods using the same base architecture. "Param." indicates the number of tunable parameters in the pre-trained encoders. "Mem." denotes the peak GPU memory footprint with batch size 64 during fine-tuning.
We highlight the best and the second results.
(M) (GB) val testA testB val testA testB val-g val-u test-u
Fully fine-tuning 151 (100%) 38.95 (100%) 80.32 82.67 75.24 63.50 68.15 55.63 66.56 67.66 67.44
Adapter [12] 3.27 (2.17%) 28.52 (73.22%) 78.02 79.89 75.23 61.35 66.34 54.21 63.18 65.26 66.65
LoRA [13] 2.37 (1.57%) 20.37 (52.30%) 77.57 78.22 73.37 61.24 66.53 53.95 64.27 67.36 66.43
AdaptFormer [15] 2.38 (1.57%) 20.37 (52.30%) 76.32 77.16 73.94 60.96 65.19 53.88 61.81 65.44 64.37
CM Adapter [40] 3.27 (2.17%) 27.19 (69.81%) 77.37 78.81 74.07 61.34 66.10 53.31 63.93 65.75 64.72
MRS-Adapter [11] 1.58 (1.05%) 20.07 (51.53%) 77.14 77.80 74.80 61.13 66.38 53.13 63.07 66.46 65.16
M2IST (Ours) 3.19 (2.11%) 15.44 (39.64%) 81.35 82.29 77.98 63.15 67.11 55.52 67.50 67.67 67.41
IV Experiments
In this section, we first introduce the datasets and evaluation metrics used to assess the performance and efficiency of our method in Section IV-A. Next, we describe the implementation details in Section IV-B. Section IV-C provides comprehensive comparisons of performance and fine-tuning efficiency, followed by an extensive ablative study and analysis of design choices in Section IV-D. Finally, we present multiple qualitative results to analyze model behavior in Section IV-E.
IV-A Datasets and Evaluation Metrics
Datasets
To verify the effectiveness and efficiency of our method, we conduct experiments on the following REC benchmarks as follows: (1) RefCOCO [22] consists of 19,994 images with 142,210 referring expressions for 50,000 objects. The RefCOCO dataset is officially split into train, validation, testA, and testB sets containing 120,624, 10,834, 5,657, and 5,095 expressions, respectively. (2) RefCOCO+ [22] includes 19,922 images with 141,564 referring expressions for 49,856 objects. Compared to RefCOCO, the referring expressions in RefCOCO+ focus more on attributes of the referred objects, such as color and shape, without including any positional words. (3) RefCOCOg [23, 24] contains 25,799 images with 95,010 referring expressions for 49,822 objects. Compared to RefCOCO and RefCOCO+, the referring expressions in RefCOCOg are typically longer, averaging almost twice the length of those in the other two datasets. RefCOCOg has two commonly used split strategies: the google split [23] (-g) and the umd split [24] (-u).
Evaluation Metrics
Following previous work [3, 25, 26], we conduct experiments on both RefCOCOg-g (val-g) and RefCOCOg-u (val-u and test-u). We use Precision@0.5 as the evaluation metric. In addition to accuracy, we also report the number of tunable parameters in the pre-trained encoders and the training memory consumption in Gigabytes (GB) to compare the fine-tuning efficiency with other PETL methods.
IV-B Implementation Details
Model Weights
The Vision Encoder is initialized with the backbone (i.e., ResNet-50 [45]) and encoder weights from DETR [20], which is pre-trained on the entire MS-COCO dataset [49]. Specifically, during the pre-training of the Vision Encoder, images from the validation and test sets of RefCOCO/+/g that overlap with MS-COCO [49] are excluded. The Language Encoder is initialized with BERT-base [21], pre-trained on the BookCorpus [50] and English Wikipedia [21]. The Vision-Language (V-L) Encoder is initialized using Xavier initialization. The proposed M3ISAs are initialized with Kaiming normal initialization.
Hyper-parameters Settings
M3ISAs are inserted into the transformer encoder layers at the same indices as those in the Vision Encoder and Language Encoder, and relevant ablation study is conducted in Table VIII. Unless otherwise specified, the bottleneck dimensions of the Visual Expert Adapter (VEA) and Language Expert Adapter (LEA) $C_{d}$ are set to 128 by default, while the interaction dimension $C_{i}$ of the Interaction Expert Adapter (IEA) is 256 by default. The scaling factor $s$ for all adapters is set to 0.1.
Training Details
For RefCOCO [22] and RefCOCOg [23, 24] datasets, the entire network is trained for 90 epochs using the AdamW optimizer, with a learning rate of $10^{-4}$ for the V-L Encoder and $10^{-5}$ for the M3ISAs. The weight decay is $10^{-4}$, and the learning rate is reduced by a factor of 10 after 60 epochs. While for RefCOCO+ [22], the network is trained for 180 epochs with the same learning rates and weight decay, but the learning rate is decreased by a factor of 10 after 120 epochs. All experiments are conducted on one A800 GPU.
IV-C Main Results
In this sub-section, we compare our M2IST under different settings to evaluate its performance and the three efficiency factors discussed in Section III-C.
TABLE II: Comparison with full fine-tuning on RefCOCO, RefCOCO+, and RefCOCOg. "RN50" and "RN101" represent ResNet-50 [45], ResNet-101 [45]. "Params." is the number and average percentage of tuned parameters in the encoders.
@@ IV-C Main Results
VC [51] VGG16 LSTM 17 (100%) - 73.33 67.44 - 58.40 53.18 62.30 - -
ParalAttn [52] VGG16 LSTM 17 (100%) - 75.31 65.52 - 61.34 50.86 58.03 - -
MAttNet [1] RN101 LSTM 47 (100%) 76.65 81.14 69.99 65.33 71.62 56.00 - 66.58 67.27
RvG-Tree [30] RN101 LSTM 47 (100%) 75.06 78.61 69.85 63.51 67.45 56.66 - 66.95 66.51
@@
FAOA [2] DarkNet-53 LSTM 43 (100%) 72.54 74.35 68.50 56.81 60.23 49.60 56.12 61.33 60.26
RCCF [53] DLA34 LSTM 18 (100%) - 81.06 71.85 - 70.35 56.32 - - 65.73
ReSC [32] DarkNet-53 BERT 152 (100%) 76.59 78.22 73.25 63.23 66.64 55.53 63.12 67.30 67.20
RealGIN [54] DarkNet-53 GRU 41 (100%) 77.25 78.70 72.10 62.78 67.17 54.21 - 62.75 62.33
TRAR [55] DarkNet-53 LSTM 43 (100%) - 79.60 71.30 - 65.10 53.50 - 63.30 62.50
TransVG [3] RN50+DETR BERT 151 (100%) 80.32 82.67 75.24 63.50 68.15 55.63 66.56 67.66 67.44
VGTR [35] RN50 LSTM 52 (100%) 78.70 82.09 73.31 63.57 69.65 55.33 62.88 65.62 65.30
PFOS [7] DarkNet-53 BERT 152 (100%) 77.37 80.43 72.87 63.74 68.54 55.84 61.46 67.08 66.35
DMRNet [56] DarkNet-53 BERT 152 (100%) 76.99 79.71 72.67 61.58 66.60 54.00 - 66.03 66.70
MRLN [4] VGG16 GRU 18 (100%) 81.39 83.65 75.03 66.33 69.75 58.05 - 65.52 65.08
D-MDETR [9] RN50+DETR BERT 143 (100%) 80.47 82.63 75.96 65.52 70.41 57.68 66.87 69.20 68.56
Ours:
M2IST ($C_{d}=8$) RN50+DETR BERT 1.91 (1.26%) 81.55 83.07 77.31 62.73 66.96 55.93 65.47 66.79 66.30
M2IST ($C_{d}=32$) RN50+DETR BERT 2.20 (1.46%) 81.03 82.34 77.54 62.48 66.73 55.70 66.22 67.42 67.83
M2IST ($C_{d}=128$) RN50+DETR BERT 3.19 (2.11%) 81.35 82.29 77.98 63.15 67.11 55.52 67.50 67.67 67.41
Comparison with PETL Methods
We first compare our proposed M2IST method with several existing parameter-efficient transfer learning (PETL) approaches, using the same base architecture and the same bottleneck dimensions for the adapter modules ($C_{d}=128$). These PETL methods include vanilla Adapter [12], LoRA [13], AdaptFormer [15], CM Adapter [40], and MRS-Adapter [11].
From Table I, we can see that all PETL methods including M2IST achieve significant parameter efficiency during fine-tuning by reducing the number of tunable parameters compared to full fine-tuning. Despite the parameter efficiency achieved by these PETL methods, they still face two major limitations in practical applications: a) All methods exhibit performance degradation compared to full fine-tuning. b) They require substantial GPU memory during fine-tuning, thus failing to deliver the efficiency that is crucial during implementation.
As for the performance, M2IST stands out as the only PETL method that can achieve performance on par with or even better than full fine-tuning, significantly outperforming other PETL methods across all three benchmarks. This highlights the effectiveness of M3ISAs in adapting pre-trained knowledge for the REC task. We can observe that all PETL methods exhibit distinct performance degradation compared to full fine-tuning on the RefCOCOg dataset. This is because RefCOCOg is a substantial dataset with sufficient data, reducing the likelihood of overfitting when fully fine-tuning the models. Even so, with the facilitation of cross-modality interaction between the encoders, M3ISAs are able to enhance the modeling of complex spatial relationships, leading to competitive performance with full fine-tuning on the RefCOCOg benchmark.
@@
Regarding fine-tuning efficiency, M2IST requires the least training memory among PETL methods. This results from the fact that gradients backpropagate through the lightweight M3ISAs rather than the heavy encoders, highlighting M2IST’s advantage in memory efficiency, as mentioned in Section III-C.
Comparison with Full Fine-tuning Methods
We further compare our M2IST with traditional full fine-tuned REC methods in Table II.
Table II demonstrates that M2IST achieves competitive performance across three benchmarks compared to most full fine-tuning methods with the fewest (only 1.91%) trainable backbone parameters. Specifically, on the three sets of RefCOCO [22], M2IST outperforms the majority of other fully fine-tuning REC methods. We can observe that MAttNet [1] and MRLN [4] achieve impressive performance in in RefCOCO+ [22]. As a Two-stage REC method, MAttNet introduces modular attention networks to separately model the subject, location, and relationship, which can more explicitly locate the referred objects by directly computing the similarity scores between the region proposals and the sentences, thus leading to enhanced performance in RefCOCO+. Similarly, the multiple relational learning modules in MRLN are also well-suited to capture these structured and compositional representations of the vision-language interactions. This allows MRLN to excel on the RefCOCO and RefCOCO+ , where the referring expressions tend to have a more well-defined structure and can be more effectively represented through the learned feature-feature, feature-task, and task-task relationships. However, as the referring expressions become longer and more complex (i.e., RefCOCOg [23, 24]), the limitations of methods that rely heavily on structured representation learning (MAttNet and MRLN) become more apparent. In contrast, transformer-based approaches (e.g., TransVG [3], D-MDETR [9]) that learn end-to-end vision-language representations show promising performance in these more challenging scenarios. Our proposed M2IST further strengthens the vision-language representation learning by combining the channel-level and token-level vision-language alignment, thus achieving impressive performance on the RefCOCOg benchmark.
In summary, Table II illustrates that M2IST achieves an optimal performance-efficiency trade-off compared to listed full fine-tuning methods, underscoring its advantage in parameter efficiency, as discussed in Section III-C.
TABLE III: Comparisons of training and inference speed. "Full" represents fully fine-tuning the baseline modal. "Training" denotes training on RefCOCO, while "Inference" denotes inference time on RefCOCO val.
Speed Full Adapter M2IST
Training (min/epoch) $\downarrow$ 52 38 33
Inference (s) $\downarrow$ 145 146 145
Comparison of Time Efficiency
We present the training and inference speeds on RefCOCO dataset of full fine-tuning, standard adapter-tuning [12], and our M2IST in Table III.
For training speed, it is evident that PETL methods improve training speed compared to full fine-tuning. This improvement is largely due to the reduced number of parameters that need to be updated when using PETL methods. Notably, M2IST achieves greater efficiency than adapter-tuning in terms of training time by updating parameters in parallel with pre-trained encoders, achieving approximately 36.54% faster than full fine-tuning. For inference speed, while adapter-tuning slightly reduces inference speed, our M2IST maintains it. This is attributed to the fact that, during inference, M2IST operates in parallel with the pre-trained encoders, maintaining inference speed. In contrast, standard adapters are inserted within the pre-trained encoders, leading to lower computational efficiency. These findings are consistent with the advantage in time efficiency proposed in Section III-C.
TABLE IV: Comparison on ReferItGame. Both the standard Adapter and our M2IST use the same base architecture.
[…]
(M) (GB) val test
Adapter 3.27 (2.17%) 28.52 (73.22%) 57.28 56.89
M2IST 3.19 (2.11%) 15.44 (39.64%) 60.61 59.30